4) For a fixed rate, a fixed principal amount, and a fixed compounding cycle, the return is an exponential function of time. Using the formula,A= P(1+ r/n)^ nt,let r = 8%, P = 1, and n = 1 and give the coordinates (t,A) for the points where t = 0, 1, 2, 3, 4.
Answer by Nate(3500)

4) For a fixed rate, a fixed principal amount, and a fixed compounding cycle, the return is an exponential function of time. Using the formula,A= P(1+ r/n)^ nt,let r = 8%, P = 1, and n = 1 and give the coordinates (t,A) for the points where t = 0, 1, 2, 3, 4.
A=(1.08)^(t)
for (t,A):
(0,1)
(1,1.08)
(2,1.1664)
(3,1.259712)
(4,1.36048896)